How do you solve and check: 8^(2x+3)=4*2^(x+1)
Answer by lwsshak3(11628)

How do you solve and check:
8^(2x+3)=4*2^(x+1)
take log of both sides
(2x+3)log8=log4+(x+1)log2
2xlog8+3log8=log4+xlog2+log2
2xlog8-xlog2=-3log8+log4+log2
x(2log8-log2)=-3log8+log4+log2
x=(-3log8+log4+log2)/(2log8-log2)
using calculator
x=-1.2
check:
8^(2x+3)=8^(-2.4+3)=8^.6≈3.4822..
4*2^(x+1)=4*2^(-1.2+1)=4*2^-.2≈3.4822
left side≈right side